if point p(x,y) is equidistant from the points (-a-b,a+b) and (a-b,a,+b) prove that x-a=0

Grade:10

1 Answers

Vijay Mukati
7 years ago
Dear Student,

Since Point P is equidistant and thus PA = PB where A(-a-b, a+b) and B(a-b, a+b)
or PA2 = PB2
i.e. (x + a +b)2 + (y-a-b)2 = (x-a+b)2 + (y-a-b)2
or (x + a + b)2 = (x – a + b)2
on further simplifying, we get
2ax = 2ab or x-b=0
(You asked wrong prrof or may be the points A and B may be incorrect.)
Thanks,
Vijay